# Quantum Vacuum energy as the origin of Gravity

### v2

@@ I Introduction
With the above introductory remarks, let us summarize the remainder of this article. The two primary assumptions we make are the formula (3) for $\rho_{\rm vac}$} which was well-motivated by the works ALrhovac1 ; ALrhovac2 , and the renormalization group flow for $\mathfrak{g}$ in equation (35) below. In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide an argument for how gravity can arise from $\rho_{\rm vac}$}. In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations. The formula (2) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$. The analysis in Section II is independent of the form $\rho_{\rm vac}$} in (3), and we turn to understanding the implications of (3) in Section III. There we assume $\mathfrak{g}$ is a marginal coupling with a renormalization group (RG) flow, which as we will argue is well-motivated, and show that this leads to an effective, induced RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$. [^11] We previously published some preliminary work on the possible implications of the RG flow of $\mathfrak{g}$ for cosmology in LeClairUniverse , however the present article supersedes it. As is standard in Cosmology, we tie this energy scale $\mu$ to the redshift $z$ and obtain equation (43). Solving the resulting equations for the scale factor $a(t)$ results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric singularity since for all times $a(t)>a_{\rm min}$. The solution to the scale factor satisfies an interesting symmetry between the far past and far future, namely $a(t)=a(-t+2t_{\rm min})$ where $a(t_{\rm min})=a_{\rm min}$. There is no Big Bang in our model since the solutions in past and future match at the time $t_{\rm min}$. Extending our solution to the far past, the time evolution of the Universe is more analogous to the swing of a pendulum. Our interpretation of this is that the age of the Universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$. In Section IV we add matter and radiation, and due to the RG flow of $G_{N}$ in (43) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold. This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}$ is small. We apply these ideas to the so-called Hubble tension, wherein the discrepancy of measurements of Hubble constant $H_{0}$ is proposed to arise from the fact that the two contradictory measurements rely on data that refers to different epochs where Newton’s constant is effectively different since the energy scales involved are different.
@@ II Emergence of Einstein’s equations from a Gravitational Casimir effect
In this section we formulate a gravitational version of the electromagnetic Casimir effect. Without assuming Einstein’s field equations, we present heuristic arguments for how they emerge in this context. These arguments lead to the formula (2) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in the Universe. This doesn’t rely on the explicit formula (3) for $\rho_{\rm vac}$}, where its implications will be studied in the next section.
@@
Let us first review the electromagnetic Casimir effect, which is well understood theoretically and has been measured Casimir . This Casimir effect requires two uncharged, parallel conducting plates in vacuum separated by a distance $L$. The energy density of quantized photons between the plates $\rho_{\rm rad}$ leads to a measurable force between the plates $F=\partial_{L}E$ where $E=\rho_{\rm rad}\cdot V$ where $V$ is the volume between the plates. It is important to point out that the theoretical predictions, which are confirmed in the measurements, do not measure the vacuum energy density $\rho_{\rm vac}$} but rather the energy density of free photons to leading order, whose pressure is $p=w\rho_{\rm rad}$, where $w=1/3$, whereas genuine vacuum energy density has w=1w=-1. In fact, since the free theory of photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$.
II.1 Formulating a gravitational Casimir effect without plates
@@
For gravity, there is no well-formulated analog of the electromagnetic Casimir effect, since hypothetical gravitons cannot be constrained by material boundaries like plates. In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable. This led us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity. Throughout this paper we assume ${\rho_{\rm vac}}>0$ such that we are dealing with de Sitter space. We make three additional assumptions: (i) $\rho_{\rm vac}$} is well-defined and finite in flat Minkowski space for the entirety of quantum fields for matter and radiation without gravity, defined in (1). This quantity $\rho_{\rm vac}$} is in principle computable without gravity, as explained in the Introduction. Secondly, (ii) The Vacuum has no time dependence originating from Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$. However as we will see in the next section, tying the RG energy scale to the redshift $z$ leads to an additional induced time dependence. Third, (iii) There is no additional and independent free parameter $\Lambda$ in Einstein’s equations corresponding to a classical cosmological constant. This is consistent with the fact that $\rho_{\rm vac}$} is not invariant under arbitrary shifts of the potential energy since it is in principle computable from the logarithm of the partition function and not its derivatives.
@@
We carry out this construction in arbitrary spatial dimension $d$ in order to make certain points. Consider first a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}. There are an infinite number of integrable QFT’s where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 . Let the spatial dimension be a finite size segment of length $2R$. Since total vacuum energy is $E=2R\,{\rho_{\rm vac}}$, there is a force $F=dE/dR=2{\rho_{\rm vac}}$ that can in principle be measured by moving the “walls”, in this case simply the endpoints of the segment. However if the endpoints do not consist of matter being held in place, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}. This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$:
@@
The dimensionless scale factor can be thought of as $a(t)=R(t)/R_{0}$ where $R_{0}$ is the scale factor today, such that at the present time $t_{0}$, $a(t_{0})=1$. In order to make sense of Newton’s 2nd law, one needs to introduce an inertial mass $M$. One also needs to introduce a length scale $R$ such that $R\ddot{a}$ is an acceleration, where as usual over-dots refer to time derivatives. Equating the force $F$ with $2{\rho_{\rm vac}}$, and requiring both sides scale the same way with the scale factor $a$, one obtains
@@
We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is $\ddot{a}/a=(\dot{a}/a)^{2}$:
@@
Let us extend the above arguments to $d$ spatial dimensions, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)}.$ (13)
The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force. To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
@@
Thus far, Newton’s constant $G_{N}$ has played no role, since we have not assumed Newtonian gravity nor Einstein’s general relativity. $G_{N}$ can be identified by comparing with Einstein’s field equations for this particular metric. We adopt standard conventions for the Einstein-Hilbert action in $d$ dimensions:
@@
The metric is as in (10), with with $dx^{2}$ replaced with $d\vec{x}\cdot d\vec{x}$. There is no a priori reason to include a curvature term proportional the standard $k\in\{-1,0,1\}$ in the spatial metric based on the above formulation of the Casimir effect. In other words $k=0$ is natural for our formulation, since, from our perspective, gravity originates from $\rho_{\rm vac}$} in Minkowski space, and the latter clearly has $k=0$. This leads to the field equations
@@
If one carefully keeps track of the $d$-dependence, this leads to the following Friedmann equations [^12] The $d$ dependence of this formula should be a well-known result. One can find it for instance in Gibbons .
@@
which is referred as the Hubble radius. With the identification of $G_{d}$ in (21), the equation (20) is equivalent to
@@
Let us for the moment express (26) as $G_{N}=3c^{2}H^{2}/8\pi{\rho_{\rm vac}}$. If one relies only on experimental data in modern cosmology, then the above expression (26) for $G_{N}$ is approximately correct with some assumptions, however it should be realized that at this stage this is a tautology since analysis of the cosmological data depends on $G_{N}$. First of all, we have not yet included matter and radiation as excitations over the Vacuum, where $H$ is no longer a constant, and we will return to this below. Let us simply point out that if one takes the experimental value (7) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (52) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$. Let us note that the same computation using $H_{0;{\rm SN}}$ in (52) yields a higher value for $G_{N}$. This will be addressed below in connection with the “Hubble tension” by taking into account the explicit expression (3) for $\rho_{\rm vac}$}.
II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
There are some important results in semi-classical gravity where the gravitational field is not quantized, namely they do not rely on the existence of gravitons, but the result still depends on the Planck length $\ell_{p}$, such as the Bekenstein-Hawking entropy and temperature for black holes Bekenstein ; Hawking . In this present context, the analog is the Gibbons-Hawking entropy for de Sitter space GibbonsHawking . Since these results depend on $G_{N}$, the expression (2) for $G_{N}$ leads to an interesting new interpretation of this entropy.
@@ II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
The Gibbons-Hawking entropy was thus proposed to be
@@
where $A=4\pi{{\rm R}^{2}_{\scriptscriptstyle\infty}}$. Here each bit is commonly interpreted as a pixelation of the area $A$ into quantized units of area $4\ell_{p}^{2}$, suggesting that spacetime itself is somehow quantized into these minimal area units at the horizon. Such a quantization of spacetime itself remains a somewhat vague notion, and difficult to make sense of mathematically.
@@
Let us point out that the above interpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole. The reason is that (2) is the correct relation between the mass $M_{\bullet}$ of the black hole and its radius $R_{\bullet}$ if one makes the replacement ${{\rm M}_{\scriptscriptstyle\infty}},{{\rm R}_{\scriptscriptstyle\infty}}\to M_{\bullet},R_{\bullet}$:
@@
From this perspective, $G_{N}$ is a universal constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,

### v3

@@ I Introduction
With the above introductory remarks, let us summarize the remainder of this article as outlined in the Table of Contents. In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide a heuristic argument for how gravity can arise from $\rho_{\rm vac}$}. In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations where $G_{N}$ is given by (6). The formula (6) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles like photons at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$, as described in Section III. The analysis in Section II is independent of the formula for $\rho_{\rm vac}$} in (2), and we turn to understanding the implications of it in Section IV. There we argue that the RG flow for the coupling $\mathfrak{g}$ induces an effective RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$. [^10] We previously published some preliminary work on the possible implications of the RG flow of $\mathfrak{g}$ for cosmology in LeClairUniverse , however the present article supersedes it. As is standard in Cosmology, we tie this energy scale $\mu$ to the redshift $z$ and obtain equation (55). In Section V we solve exactly the modified Friedmann equations for the dark-universe, which results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric curvature singularity since for all times $a(t)>a_{\rm min}$. The solution to the scale factor satisfies an interesting symmetry between the far past and far future, namely $a(t)=a(-t+2t_{\rm min})$ where $a(t_{\rm min})=a_{\rm min}$. There is no Big Bang normally associated with the singularity $a(t)=0$, rather at the time $t_{\rm min}$ the universe is at its hottest, and is more properly referred to as a hot Big Bang. In our model since the solutions in past and future match at the time $t_{\rm min}$, one can address what happened before the Big Bang. Extending our solution to the far past, the time evolution of our model universe is more analogous to the harmonic swing of a pendulum. Our interpretation of this is that the age of this model universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$. In Section VI we add matter and radiation, and due to the RG flow of $G_{N}$ in (55) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold. This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}\equiv\,b\,\mathfrak{g}_{0}/2\pi$ is small. We apply these ideas to the so-called Hubble tension, wherein the discrepancy of measurements of Hubble constant $H_{0}$ is proposed to arise from the fact that the two contradictory measurements rely on data that refers to different epochs where Newton’s constant is effectively different since the energy scales involved are different. In Section VII we perform some checks of our model by comparing with a variety of cosmological and other types of experimental data. The strongest support comes from recently observed trends in the Hubble tension for low-z supernovae.
@@ II Emergence of Einstein’s equations from a Gravitational Casimir effect
In this section we formulate a gravitational version of the electromagnetic Casimir effect. These considerations are independent of the specific QFT details that lead to the formula (2). In this section we only assume that it is well-defined and finite, and will turn to the implications of the formula (2) in the next section. Without assuming Einstein’s field equations, we present heuristic arguments for how they emerge in this context. These arguments lead to the formula (6) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in our model universe. The discussion of this section is not absolutely necessary for understanding the remainder of this article. The reader may prefer to simply assume Einstein’s field equations and move onto the next section where QFT properties of $\rho_{\rm vac}$} are invoked.
@@
Let us first review the electromagnetic Casimir effect, which is well understood theoretically and has been measured Casimir . This Casimir effect requires two uncharged, parallel conducting plates in vacuum separated by a distance $L$. The plates are required to radiate and confine photons between the plates. The energy density of quantized photons between the plates $\rho_{\rm rad}$ leads to a measurable force between the plates $F=\partial_{L}E$ where $E=\rho_{\rm rad}\cdot V$ where $V$ is the volume between the plates. It is important to point out that the theoretical predictions, which are confirmed in the measurements, do not measure the vacuum energy density $\rho_{\rm vac}$} but rather the energy density of free photons to leading order, whose pressure is $p=w\rho_{\rm rad}$, where $w=1/3$, whereas genuine vacuum energy density has w=1w=-1. In fact, since the theory of non-interacting photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$ in the leading conformal limit.
@@
For Gravity, there is no well-formulated and strict analog of the electromagnetic Casimir effect, since gravitational waves and hypothetical gravitons cannot be constrained by material boundaries like conducting plates. In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable if they originate from gravitons. This leads us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity. Throughout this paper we assume ${\rho_{\rm vac}}>0$ such that we are dealing with de Sitter space. Our working assumptions were presented in the Introduction. One additional assumption is that the vacuum has no time dependence originating directly from the time dependence of the resulting Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$. However as we will see in the next section, tying the RG energy scale to the redshift $z$ leads to an additional induced time dependence.
@@
We carry out this construction in arbitrary spatial dimension $d$ in order to make certain conceptual points. Let us begin with only $1$ more spatial dimension than the harmonic oscillator in QM discussed in the Introduction which corresponds to $d=0$. Thus, consider a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}. There are an infinite number of integrable QFT’s for $d=1$ where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 which are reviewed in Section IVA. Let the spatial dimension be a finite size segment of length $2R$. Since total vacuum energy is $E=2R\,{\rho_{\rm vac}}$, there is a force $F=dE/dR=2{\rho_{\rm vac}}$ that can in principle be measured by moving the “walls”, in this case simply the endpoints of the segment. However if the endpoints do not consist of matter being held in place, that is to say there is no analog of conducting plates, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}. This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$ which can incorporate this spatial expansion:
@@
The hypothetical experimental setup is the Universe itself, since there are no conducting plates involved, and as we will argue, the cosmic horizon plays the role of such plates. The dimensionless scale factor by definition satisfies $a(t_{0})=1$ at the present time $t_{0}$.
In order to make sense of Newton’s 2nd law, one needs to introduce an inertial mass $M$. One also needs to introduce a length scale $R$ such that $R\ddot{a}$ is an acceleration, where as usual over-dots refer to time derivatives. Equating the force $F$ with $2{\rho_{\rm vac}}$, and requiring both sides scale the same way with the scale factor $a$, one obtains
@@
We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is just $\ddot{a}/a=(\dot{a}/a)^{2}$:
@@
Let us extend the above arguments to $d$ spatial dimensions, and assume rotational symmetry, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)},$ (11)
where $\Gamma$ is the usual $\Gamma$-function. The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force. To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
@@
Thus far, Newton’s constant $G_{N}$ has played no role, since we have not assumed Newtonian gravity nor Einstein’s general relativity. $G_{N}$ can be identified by comparing with Einstein’s field equations for this particular metric. We adopt standard conventions for the Einstein-Hilbert action in $d$ spatial dimensions:
@@
The metric is as in (8), with with $dx^{2}$ replaced with $d\vec{x}\cdot d\vec{x}$. Based on the above formulation of the gravitational Casimir effect, there is no a priori reason to include a curvature term in the spatial part of the metric proportional the standard $k\in\{-1,0,1\}$. In other words $k=0$ is natural for our formulation, since, from our perspective, Gravity originates from $\rho_{\rm vac}$} in Minkowski space, and the latter clearly has $k=0$. This leads to the field equations
@@
It’s important to note that there is no room, i.e. no a priori reason, to introduce an additional classical free parameter $\Lambda$ corresponding to an independent cosmological constant term in Einstein’s equations based on the above arguments. Furthermore, we already explained that once the QFT is decided upon, there is no room for arbitrary shifts in $\rho_{\rm vac}$} since it is based on the logarithm of its free energy density and not its derivatives, and this avoids the fine-tuning issues normally associated with a variable free parameter $\Lambda$ corresponding to a cosmological constant. If one carefully keeps track of the $d$-dependence, this leads to the following Friedmann equations [^11] The $d$ dependence of this formula should be a well-known result. One can find it for instance in GibbonsDdim .
@@
which is referred as the Hubble radius. With the identification of $G_{d}$ in (19), the equation (18) is equivalent to equation (14):
@@
Let us for the moment express (24) as $G_{N}=3c^{2}H^{2}/8\pi{\rho_{\rm vac}}$. If one relies only on experimental data in modern cosmology, then the above expression (24) for $G_{N}$ is approximately correct with some assumptions, however it should be realized that at this stage this is a tautology since analysis of the cosmological data depends on $G_{N}$. First of all, we have not yet included matter and radiation as excitations over the Vacuum, where $H$ is no longer a constant, and we will return to this below. Let us just point out that if one takes the experimental value (5) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (64) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$. Let us note that the same computation using $H_{0;{\rm SN}}$ in (64) yields a higher value for $G_{N}$. This will be addressed below in connection with the “Hubble tension” by taking into account the explicit expression (2) for $\rho_{\rm vac}$}.
III A reinterpretation of the entropy of de Sitter space and black holes
There are some important results in semi-classical gravity where the gravitational field is not quantized, namely they do not rely on the existence of gravitons, but the result still depends on the Planck length $\ell_{p}$, such as the Bekenstein-Hawking entropy and temperature for black holes Bekenstein ; Hawking . In this present context, the analog is the Gibbons-Hawking entropy for de Sitter space GibbonsHawking . Since we have brought into scrutiny whether $G_{N}$ and thus Planck length $\ell_{p}$ are fundamental parameters, this led us to re-interpret these entropies, since they are expressed in terms of $\ell_{p}$. The expression (6) for $G_{N}$ leads to such a new interpretation that is very simple and illuminating.
@@ III A reinterpretation of the entropy of de Sitter space and black holes
The Gibbons-Hawking entropy $S_{GH}$ was proposed to be
@@
where $A=4\pi{{\rm R}^{2}_{\scriptscriptstyle\infty}}$. Here each bit is commonly interpreted as a pixelation of the area $A$ into quantized units of area $4\ell_{p}^{2}$, suggesting that spacetime itself is somehow quantized into these minimal area units at the horizon. The overall factor of $1/4$ is the hardest to explain in the above formula. Such a quantization of spacetime itself remains a somewhat vague notion, and difficult to make sense of mathematically.
@@
The above reinterpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole, where again the bits are quantized massless particles. [^14] For super massive black holes, the quantization of spacetime itself into bits at the super large horizon seems to us as rather implausible. The reason is that (6) is the correct relation between the mass $M_{\bullet}$ of the black hole and its radius $R_{\bullet}$ if one makes the replacement ${{\rm M}_{\scriptscriptstyle\infty}},{{\rm R}_{\scriptscriptstyle\infty}}\to M_{\bullet},R_{\bullet}$:
@@
From this perspective, $G_{N}$ is a constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,